# Lost-in-Distance: Impact of Contextual Proximity on LLM Performance in Graph Tasks

### v1

@@ Abstract
Despite significant advancements, Large Language Models (LLMs) exhibit blind spots that impair their ability to retrieve and process relevant contextual data effectively. We demonstrate that LLM performance in graph tasks with complexities beyond the “needle-in-a-haystack” scenario—where solving the problem requires cross-referencing and reasoning across multiple subproblems jointly—is influenced by the proximity of relevant information within the context, a phenomenon we term “lost-in-distance”. We examine two fundamental graph tasks: identifying common connections between two nodes and assessing similarity among three nodes, and show that the model’s performance in these tasks significantly depends on the relative positioning of common edges. We evaluate three publicly available LLMs—Llama-3-8B, Llama-3-70B, and GPT-4—using various graph encoding techniques that represent graph structures for LLM input. We propose a formulation for the lost-in-distance phenomenon and demonstrate that lost-in-distance and lost-in-the middle phenomenas occur independently. Results indicate that model accuracy can decline by up to 6x as the distance between node connections increases, independent of graph encoding and model size.
@@ 1 Introduction
Particularly, by leveraging the extensive common knowledge and powerful semantic comprehension abilities of LLMs, recent research has aimed to apply them to tasks related to graph structures (Wang et al., 2024). LLMs are increasingly being adopted for a variety of tasks that involve graph structures, such as planning in robotics (Andreas, 2022), knowledge extraction using knowledge graphs (Shen et al., 2020; Saxena et al., 2020), and multi-hop question answering (Creswell et al., 2022; Fang et al., 2019). For instance, they have been used to guide agents through structured graph-based environments (Huang et al., 2022). Building upon these applications, recent works by Sanford et al. (2024), Perozzi et al. (2024), and Agarwal et al. (2020) have demonstrated that graph tasks can be encoded into textual formats that allow pre-trained LLMs to solve them as out-of-domain tasks. This innovative approach effectively transforms graph problems into a language that LLMs can understand and process.
@@
While LLMs are being expanded in many applications, they suffer from certain blind spots that significantly affect their performance. In particular, how these models process information in their context and retrieve relevant data to solve the task at hand remains an active area of research (Kaddour et al., 2023). Understanding these limitations is crucial for extending context length (Gemini et al., 2023; Xu et al., 2023; Chen et al., 2023) and improving in-context learning (Zhou et al., 2022; Wei et al., 2023; An et al., 2024). Recent works have shown that the performance of LLMs depends on the location of information in their context. Primarily, Liu et al. (2023) introduces the “lost-in-the-middle” phenomenon, where information placed in the middle of a prompt is less effectively utilized by the model compared to information at the beginning or end, resulting in significant performance degradation when the position of relevant information in the context changes.
Unlike these previous works which mainly focus on shortcomings of LLMs in NLP tasks, e.g. lost-in-the-middle, we focus on the deficiency of these models in tasks beyond natural language processing, specifically solving fundamental graph problems. This area is heavily under-explored and has wide practical applications (Perozzi et al., 2024; Colon-Hernandez et al., 2021; Xie et al., 2023). Since these tasks require understanding graph structure and relationship between objects, they provide us with great insights into model’s blind spots. Through our analysis, we provide insight that LLMs not only have blind spots regarding where information exists in the context, but their performance in solving complex tasks also depends on the relative position of information within the context.
@@
Particularly, we look into Common Connection and Similarity tasks, which are the main algorithms used as the backbone of many applications such as molecular design (Tan et al., 2023; Xia et al., 2023), social network analysis (Gao et al., 2024), and recommendation systems (Li et al., 2023). For example, these tasks are the main algorithms in “user-user” and “user-item” recommendations in large industry recommendation products (Xie et al., 2023; Huang et al., 2015; Wu et al., 2022; Dai et al., 2024). These tasks not only require understanding of subgraph structures but also demand integration of information and reasoning across subgraphs. We demonstrate that strong, publicly available LLMs universally degrade in performance as when relevant pieces of information are distant from each other. Our analysis shows that this effect is present even when one controls for the effects absolute position of the relevant information in the context. To summarize
@@ 3 Lost-in-the-middle for Edge Existence
    Randomly select nine additional nodes as noise nodes and incorporate their textual subgraph encodings into the prompt. This step is necessary to examine the impact of the position of relevant information.
@@ 3.1 Experimental Results
Lost-in-the-middle can happen in the edge existence task. To demonstrate the lost-in-the-middle phenomena in edge existence task, we experiment with the state of the art model as of writing this paper GPT-4. The experiment results are averaged over twenty randomly generated graph where from each graph we randomly select two nodes and form the edge existence prompt as described in previous section.
Figure 3 shows that all encodings can cause the LLM to lose the information in the middle of the prompt. The best performance occurs when the relevant information is either at the beginning or the end of the entire subgraph structure. Even for the incident encoding which has the best performance among all encodings, the LLM still has the worst performance when the answer is located in the middle of the prompt.
@@ 5.1 Experimental Setup
Leveraging in-context learning (Dong et al., 2022; Wei et al., 2023), we conducted experiments using both closed-source models (GPT-4) and open-source models (Llama-3-8B-Instruct and Llama-3-70B-Instruct). For all models, we set the decoding temperature to zero to ensure the generation of deterministic answers. In each sample, we randomly selected two or three nodes as the nodes of interest for the common connection and similarity tasks, respectively. We performed experiments on hundreds of thousands of randomly generated graphs to draw statistically significant conclusions regarding LLM behavior. The experimental results were then averaged across multiple samples.
@@ 5.3 Similarity Task
An example of the prompt for the similarity task is presented in Appendix C.1. Since the similarity task inherently involves solving two common connection tasks and a comparison task, we adopt Chain-of-Thought prompting (Wei et al., 2022) to guide the model in solving the task step by step, thereby obtaining the most accurate answers. Detailed descriptions of the prompts and further analysis of the LLMs’ failure rates in following CoT instructions, along with examples of answer degeneration, are provided in Appendix A.2. GPT-4 and Llama-3-70B exhibit low failure rates, while Llama-3-8B demonstrates a high failure rate across all graph encodings, including a 69% failure rate in expert encodings.
@@ A.2 Answer degeneration
  • Self-contradiction: In CoT prompting, LLMs are asked to answer the main question based on their responses to subquestions. However, we find that LLMs sometimes provide an incorrect final conclusion. For example, as shown in Figure 9 where we ask the LLM “is the number of common connections between node 658 and node 535 greater than the number of common connections between node 535 and node 807?”, the LLM determines that the number of common connections between node $658$ and node $535$ is $6$, and between node $535$ and node $807$ is $4$, but the final answer is “no” when it should be “yes”.
@@ C.1 Prompt Example
Figure 11: Example of the CoT prompting for the similarity task.
@@ C.2 All results
Figure 12 presents the results of the similarity task at a density of 0.1 across three models (GPT-4o, Llama-8B, Llama-70B) and three different graph encodings. For all models utilizing the graph encoding functions, we observe the typical lost-in-distance pattern, where performance at the $(\mathrm{Small},\mathrm{Small})$ index is better than at the $(\mathrm{Large},\mathrm{Large})$ index.

### v2

@@ Abstract
Despite significant advancements, Large Language Models (LLMs) exhibit blind spots that impair their ability to retrieve and process relevant contextual data effectively. We demonstrate that LLM performance in graph tasks with complexities beyond the “needle-in-a-haystack” scenario—where solving the problem requires cross-referencing and reasoning across multiple subproblems jointly—is influenced by the proximity of relevant information within the context, a phenomenon we term “lost-in-distance”. We examine two fundamental graph tasks: identifying common connections between two nodes and assessing similarity among three nodes, and show that the model’s performance in these tasks significantly depends on the relative positioning of common edges. We evaluate three publicly available LLMs using various graph encoding techniques that represent graph structures for LLM input. We propose a formulation for the lost-in-distance phenomenon and demonstrate that lost-in-distance and lost-in-the-middle phenomena occur independently. Results indicate that model accuracy can decline by up to 6x as the distance between node connections increases, independent of graph encoding and model size.
@@ 1 Introduction
Particularly, by leveraging the extensive common knowledge and powerful semantic comprehension abilities of LLMs, recent research has aimed to apply them to tasks related to graph structures (Wang et al., 2024). LLMs are increasingly being adopted for a variety of tasks that involve graph structures, such as planning in robotics (Andreas, 2022; Huang et al., 2022), knowledge extraction using knowledge graphs (Shen et al., 2020; Saxena et al., 2020; Sun et al., 2023), and multi-hop question answering (Creswell et al., 2022; Fang et al., 2019).
Building upon these applications, Chen et al. (2024) and Luo et al. (2024) introduce GraphWiz and GraphLM, respectively—fine-tuned LLMs on graph instruction and reasoning tasks—which exhibit superior performance compared to general large language models in graph understanding and reasoning capabilities. Recent works by Sanford et al. (2024), Perozzi et al. (2024), and Agarwal et al. (2020) have demonstrated that graph tasks can be encoded into textual formats, enabling general pre-trained LLMs to solve them as out-of-domain tasks. This innovative approach effectively transforms graph problems into a language that LLMs can understand and process.
@@
While general LLMs are being expanded in many applications including graph understanding, they suffer from certain blind spots that significantly affect their performance. In particular, how these models process information in their prompt and retrieve relevant data to solve the task at hand remains an active area of research (Kaddour et al., 2023). Understanding these limitations is crucial for extending context length (Gemini et al., 2023; Xu et al., 2023; Chen et al., 2023) and improving in-context learning (Zhou et al., 2022; Wei et al., 2023; An et al., 2024). Recent works have shown that the performance of LLMs depends on the location of information in their prompt. Primarily, Liu et al. (2023) introduces the “lost-in-the-middle” phenomenon, where information placed in the middle of a prompt is less effectively utilized by the model compared to information at the beginning or end, resulting in significant performance degradation when the position of relevant information in the context changes.
Unlike previous works that primarily focus on the limitations of LLMs in natural language processing tasks, this study investigates the deficiencies of these models in tasks beyond NLP, focusing on solving fundamental graph problems. While leveraging LLMs to address graph-related problems has broad practical applications (Perozzi et al., 2024; Colon-Hernandez et al., 2021; Xie et al., 2023), recent studies have shown that the sequential order in which a graph is described significantly affects LLMs’ comprehension of graph structures and their performance in solving graph tasks (Ge et al., 2024). In this study, we aim to explore the underlying phenomena and blind spots that contribute to performance variability based on the ordered description of graphs. Our analysis reveals that LLMs not only have blind spots regarding the locations of information within the prompt but that their ability to solve complex tasks also depends on the relative positioning of this information. Since graph tasks require understanding graph structure and the relationships between nodes and edges, the sequential arrangement of data plays a crucial role in how well LLMs can integrate and reason over that information. These insights suggest that improvements in graph comprehension by LLMs must account for both the placement and relative distance of information to enhance performance in graph-related problem-solving.
@@
Particularly, we look into Common Connection and Similarity tasks, which are the main algorithms used as the backbone of many applications such as molecular design (Tan et al., 2023; Xia et al., 2023), social network analysis (Gao et al., 2024), and recommendation systems (Li et al., 2023). For example, these tasks are the main algorithms in “user-user” and “user-item” recommendations in large industry recommendation products (Xie et al., 2023; Huang et al., 2015; Wu et al., 2022; Dai et al., 2024). These tasks not only require understanding of subgraph structures but also demand integration of information and reasoning across subgraphs. We demonstrate that strong, publicly available LLMs universally degrade in performance as when relevant pieces of information are distant from each other. Our analysis shows that this effect is present even when one controls for the effects absolute position of the relevant information in the prompt. To summarize
@@
  • We formulate the goodness of fit for the lost-in-distance phenomenon and demonstrate its compounded impact when combined with the lost-in-the-middle phenomenon in solving graph-related tasks.
@@ 3 Lost-in-the-middle for Edge Existence
    Randomly select nine additional subgraphs and incorporate their textual encodings into the prompt. This step is necessary to examine the impact of the position of relevant information.
@@ 3.1 Experimental Results
Lost-in-the-middle can happen in the edge existence task. To demonstrate the lost-in-the-middle phenomena in edge existence task, we experiment with one of the state of the art model as of writing this paper GPT-4 that does not use function calling. Since the lost-in-the-middle phenomenon has been extensively studied in the literature, particularly in the context of language tasks, we limit our experimentation only to GPT-4 and use its result later to study the ”lost-in-distance” phenomenon.
Figure 3 shows that all encodings can cause the LLM to lose the information in the middle of the prompt. The experiment results are averaged over twenty randomly generated graph where from each graph we randomly select two nodes and form the edge existence prompt as described in previous section.
The best performance occurs when the relevant information is either at the beginning or the end of the entire subgraph structure. Even for the incident encoding which has the best performance among all encodings, the LLM still has the worst performance when the answer is located in the middle of the prompt.
@@ 5.1 Experimental Setup
Leveraging in-context learning (Dong et al., 2022; Wei et al., 2023), we conducted experiments using both closed-source models (GPT-4) and open-source models (Llama-3-8B-Instruct and Llama-3-70B-Instruct). For all models, we set the decoding temperature to zero to ensure the generation of deterministic answers. In each sample, we randomly selected two or three nodes as the nodes of interest for the common connection and similarity tasks, respectively. We performed experiments on We conducted experiments on one hundred thousand randomly generated graphs to derive statistically significant insights into LLM behavior. The experimental results were then averaged across multiple samples.
@@ 5.3 Similarity Task
An example of the prompt for the similarity task is presented in Appendix C.1. Since the similarity task inherently involves solving two common connection tasks and a comparison task, we adopt zero-shot step-by-step reasoning prompting (Kojima et al., 2022) to guide the model in solving the task step by step, thereby obtaining the most accurate answers. Detailed descriptions of the prompts and further analysis of the LLMs’ failure rates in following zero-shot reasoning instructions, along with examples of answer degeneration, are provided in Appendix A.2. GPT-4 and Llama-3-70B exhibit low failure rates, while Llama-3-8B demonstrates a high failure rate across all graph encodings, including a 69% failure rate in expert encodings.
@@ 6 Goodness of Fit for Lost-in-Distance
7 Conclusion
This study introduces the ”lost-in-distance” phenomenon, demonstrating how the performance of LLMs in graph-related tasks is impacted by the relative positioning of relevant information in the input context. Our experiments reveal that as the distance between crucial data points increases, the accuracy of publicly available open and proprietary LLMs significantly deteriorates in solving graph tasks that require cross-referencing. This phenomenon compounds with the known ”lost-in-the-middle” effect, leading to notable declines in task performance when multiple subgraphs must be cross-referenced, such as in common connection and similarity tasks.
Our findings, which we coin ”lost-in-distance,” highlight a key limitation of current LLMs in solving graph tasks, particularly those requiring complex reasoning over dispersed data. This study paves the way for future research focused on mitigating these issues, such as developing advanced graph encoding techniques. Addressing the challenges posed by lost-in-distance will be critical for enhancing LLM applications in domains like recommendation systems, molecular design, and multi-hop question answering, and will contribute to the broader field of LLMs in graph problem-solving.
@@ A.2 Answer degeneration
  • Self-contradiction: In zero-shot step-by-step reasoning prompting (Kojima et al., 2022), LLMs are asked to answer the main question based on their responses to subquestions. However, we find that LLMs sometimes provide an incorrect final conclusion. For example, as shown in Figure 9 where we ask the LLM “is the number of common connections between node 658 and node 535 greater than the number of common connections between node 535 and node 807?”, the LLM determines that the number of common connections between node $658$ and node $535$ is $6$, and between node $535$ and node $807$ is $4$, but the final answer is “no” when it should be “yes”.
@@ C.1 Prompt Example
Figure 11: Example of the zero-shot step-by-step reasoning prompting for the similarity task.
@@ C.2 All results
Figure 12 presents the results of the similarity task at a density of 0.1 across three models (GPT-4, Llama-8B, Llama-70B) and three different graph encodings. For all models using the graph encoding functions, we observe the typical lost-in-distance pattern, where performance at the $(\mathrm{Large},\mathrm{Large})$ index—indicating a large distance between relevant information within two subgraphs in the prompt—is worse than at the $(\mathrm{Small},\mathrm{Small})$ index, which indicates a smaller distance between relevant information for the similarity task in the prompt.